OpenStudy (erinweeks):

Given: line l is parallel to line m. If the measure of angle 1 = 7x – 10, and the measure of angle 5 = 5x + 16, find the value of x and the measures of angles 1 and 5.I NEED HELP SO BAD PLEASE WALK ME THROUGH THIS ANYONE

OpenStudy (erinweeks):

@jim_thompson5910

OpenStudy (erinweeks):



OpenStudy (erinweeks):

each of the numbers are on the opposite side of the transversal from each other

OpenStudy (campbell_st):

well angle 1 and angle 5 are equal as they are corresponding angles in parallel lines. so you can say 7x - 10 = 5x + 16 solve the equation for x

OpenStudy (erinweeks):

now I did all that in my assignment and the feedback I got on it was I was also supposed to find angles 1 and 5?

jimthompson5910 (jim_thompson5910):

what did you get for x

OpenStudy (erinweeks):

x=13

OpenStudy (erinweeks):

7x-10=5x+16 +10 +10 7x=5x+26 -5x -5x 2x=26 x=13

jimthompson5910 (jim_thompson5910):

if angle 1 = 7x – 10, and x = 13, then angle 1 =???

OpenStudy (erinweeks):

81?

OpenStudy (erinweeks):

and angle 5 is parallel which means its 81 as well?

jimthompson5910 (jim_thompson5910):

yep, you got it

jimthompson5910 (jim_thompson5910):

angle 1 = angle 5 because they are corresponding angles (and because line m and l are parallel)

jimthompson5910 (jim_thompson5910):

so angle1 = 81 degrees angle5 = 81 degrees

OpenStudy (erinweeks):

Ok great I might have a few more questions for you Jim thank you :) is that ok?

jimthompson5910 (jim_thompson5910):

sure I can help, but just with a few

OpenStudy (erinweeks):

Ok thank you I only have a few so let me get those on her in one second while I finish writing my answer out on my oral component

OpenStudy (anonymous):

<1=<4 (alternate angles) <4=<5 therefore <1=<5 7x-10=5x+16 7x-5x=16+10 2x=26 x=13 <1=7*13-10=91-10=81 <1+<2=180 (adjacent angles) 81+<2=180 <2=180-81=99 <1=<4=<5=<8=81 (alternating angles) <2=<3=<6=<8=91 (corresponding angles)

OpenStudy (anonymous):

corrections <2=<3=<6=<7